# FlightScope: A Deep Comprehensive Review of Aircraft Detection Algorithms in Satellite Imagery

### v3

@@ 5.3 Results
This subsection presents a detailed result showcase of the object detection algorithms training during the 500 epochs. Fig. 11 display the mAP (overall mean avec precision accros different confidence thresholds) and the mAP50 (for object detected by an Intersection over Union (IoU) threshold of 50% and up) curves collectively visualizing the performance of the object detection algorithms in the task of aircraft detection from remote sensing imagery. Notably, the performances of the algorithms are quite comparable with an overall mAP varying between 0.86 and 0.99. Among the models, YOLOv5 emerges as a standout performer, achieving the highest mAP value of 0.99471 at step 150, showcasing its great precision and robustness. YOLOv8 closely follows, reaching a peak mAP value of 0.99236 at step 395, emphasizing the efficacy of the YOLO architecture in aerial object detection. However, SSD lags behind with a comparatively modest mAP value of 0.86 at step 74 which stabilizes in the rest of the training process.
Figure 11: Comparison of bounding box mean average precision (mAP) curves for trained object detection algorithms. To the left raw figures of the curves, the right figures are magnifications from epoch 450 to 500. (a) Represents the mAP. (b) Illustrates the mAP50
@@
In addition, Fig. 12(a) and 12(b) provide successively a visualization of the bounding box and total loss curves for the algorithms, where it is noticeable that the majority of the algorithm converge quickly around 100 epochs (apart from RTMDet which has a jump value around the epoch 280) and reach a seemingly horizontal asymptote line within 300 epochs. While extending the training duration beyond 500 epochs might promote further convergence, there is a potential risk of overfitting the models.
@@ 5.4 Evaluation on GDIT Dataset
To evaluate the efficacy of object detection models, originally trained on the HRPlanesv2 dataset, a thorough assessment was conducted using the GDIT dataset. The evaluation encompassed all subsets of images: train, test, and validation, each representing diverse scenarios. This comprehensive evaluation aimed to gauge the adaptability and robustness of the algorithms under varied conditions encountered in different subsets. A sample of unseen images from both datasets highlighting the bounding boxes of estimated aircrafts is presented in Fig. 13 and Fig. 14 where the ND, FP and IE stands respectively for ’Not Detected’, ’False Positives’ and ’Inaccurate Estimation’. The figures shows the struggle of CenterNet, Faster RCNN and SSD in the detection of small object is observable while both YOLO versions and RTMDet are able to detect over 80% of the aircraft in the image with a minimum confidence of 32.6% with small amount of FP and/or ND.
Figure 13: Inference examples of YOLOv8, DETR, SSD and CenterNet on unseen images from Google Earth HRPlanesv2 dataset and Airbus GDIT. ‘FP’ stands for False Positives, ‘ND’ for No Detection and ‘IE’ for Inaccurate Estimation.
Figure 14: Inference examples of YOLOv5, RTMDet, RetinaNet and Faster-RCNN on other set of unseen images from Google Earth HRPlanesv2 dataset and Airbus GDIT. ‘FP’ stands for False Positives, ‘ND’ for No Detection.
The results of the metrics IoU, Recall and AP are presented in the histograms (Fig. 15). On the train subset, YOLOv8 demonstrated notable performance with an AP of 91.9%, Recall of 68.6%, and IoU of 71.1%. YOLOv5 exhibited excellence with an AP of 96.8%, Recall of 66.1%, and IoU of 74.0%. Conversely, SSD displayed a comparatively lower AP of 59.4%, Recall of 38.6%, and IoU of 49.2%. Performance metrics such as AP, Recall, and IoU showed variations for Faster RCNN and CenterNet.
@@
Figure 15: Estimated evaluation metrics when inferencing the 8 models (initially trained on HRPlanesV2) on all images from unseen subsets ’Train’ (a), ’Test’ (b) and ’Validation’ (c) from GDIT aircraft dataset.
Table 4: Object Detection Performance on Remote Sensing Images. Tr. subset refers to all images from the ’train’, Te. subset refers to all images from the ’test’, Val. subset refers to all images from the ’validation’ and All dataset overall images in the dataset.
Architecture Metric Tr. subset Te. subset Val. subset All dataset
YOLOv8 AP 0.919 0.903 0.900 0.907
@@ 6 Conclusion
This study presents a comprehensive evaluation of aircraft detection algorithms in satellite imagery, namely YOLO (v5 and v8), Faster RCNN, CenterNet, RetinaNet, RTMDet and DETR, specifically focusing on the HRPlanesV2 dataset and extending the assessment to subsets of the GDIT dataset: train, test, and validation. The used training setup consists of training the aformentioned object detection algorithms on the dataset for 500 epochs on a 3 NVIDIA RTX A6000 GPUs configuration, each GPU has 48 GB of memory. The results of the evaluation demonstrate the adaptability and robustness of the trained object detection algorithms. Among these algorithms, YOLOv5 emerges as the standout performer, achieving the highest mean average precision (mAP) of 0.99, highlighting its precision and robustness. YOLOv8 closely follows, further emphasizing the effectiveness of the YOLO architecture in aerial object detection. On the other side, the SSD displayed the lowest performances in both the training and evaluation. Furthermore, the evaluation extends to the GDIT dataset, providing a more comprehensive assessment by deploying the trained network on other scenarios, including different satellite imagery sources. By employing evaluation metrics such as Average Precision, Recall, and Intersection over Union, YOLOv5 still consistently outperforms the other algorithms, demonstrating superior performance across all subsets. This solidifies YOLOv5’s position as the top-performing algorithm for aircraft detection, characterized by commendable AP, Recall, and IoU scores.
The topic of this research offers a extensive overview and comparative study, providing an in-depth analysis of the performance, accuracy, and computational demands of object detection method. The insights gained from this study significantly enhance the decision-making process for selecting the most effective aircraft localization techniques in satellite imagery, supported by detailed training and validation processes using the HRPlanesv2 dataset and additional validation with the GDIT dataset.

### v4

@@ 5.3 Results
This subsection presents a detailed result showcase of the object detection algorithm training during the 500 epochs. Figure 12 displays the mAP (overall mean average precision across different confidence thresholds) and the mAP50 (for objects detected by an intersection over union (IoU) threshold of 50% and up) curves collectively visualizing the performance of the object detection algorithms in the task of aircraft detection from remote sensing imagery. Notably, the performances of the algorithms are quite comparable with an overall mAP varying between 0.86 and 0.99. Among the models, YOLOv5 emerges as a standout performer, achieving the highest mAP value of 0.99471 at step 150, showcasing its great precision and robustness. YOLOv8 and YOLOv10 closely follow, both reaching a peak mAP value of about 0.98, emphasizing the efficacy of the YOLO architecture in aerial object detection. However, SSD lags behind with a comparatively modest mAP value of 0.86 at step 74, which stabilizes in the rest of the training process.
The results in Figure 12b confirm the previous discussion, as YOLOv5 continues to outperform, achieving the highest mAP50 value of 0.84454 at step 493. RTMDet and YOLOv8 remain strong with mAP50 values of 0.838 at step 340 and 0.8372 at step 492, respectively. CenterNet consistently performs well, achieving an mAP50 value of 0.826 at step 439. Faster RCNN maintains a balanced mAP50 value of 0.775 at step 402, showcasing reliability in detection. DETR contributes robustly with an mAP50 value of 0.774 at step 472, while RetinaNet exhibits stability with an mAP50 value of 0.765 at step 250. Finally, SSD, with a noticeable gap between mAP and mAP50 values, suggests potential challenges in localization precision, emphasizing the need for refinement in specific scenarios.
Figure 12: Comparison of bounding box mean average precision (mAP) curves for trained object detection algorithms. Raw figures of the curves are on the left; the right figures are magnifications from epochs 450 to 500. (a) represents the mAP. (b) illustrates the mAP50.
@@
In addition, Fig. 13(a) and 13(b) provide successively a visualization of the bounding box and total loss curves for the algorithms, where it is noticeable that the majority of the algorithm converge quickly around 100 epochs (apart from RTMDet which has a jump value around the epoch 280) and reach a seemingly horizontal asymptote line within 300 epochs. While extending the training duration beyond 500 epochs might promote further convergence, there is a potential risk of overfitting the models.
@@ 5.4 Evaluation on GDIT Dataset
To evaluate the efficacy of object detection models, originally trained on the HRPlanesv2 dataset, a thorough assessment was conducted using the GDIT dataset. The evaluation encompassed all subsets of images: train, test, and validation, each representing diverse scenarios. This comprehensive evaluation aimed to gauge the adaptability and robustness of the algorithms under varied conditions encountered in different subsets. A sample of unseen images from both datasets highlighting the bounding boxes of estimated aircraft is presented in Figures 14 and 15 where ND, FP, and IE stand for ‘Not Detected’, ‘False Positives’, and ‘Inaccurate Estimation’, respectively. The figures show the struggle of CenterNet, Faster RCNN, and SSD in the detection of small objects, while both YOLO versions and RTMDet are able to detect over 80% of the aircraft in the image with a minimum confidence of 32.6% with small amounts of FPs and/or ND.
The results of the metrics IoU, recall, and AP are presented in histograms (Figure 16). On the training subset, YOLOv8 demonstrated notable performance with an AP of 91.9%, recall of 68.6%, and IoU of 71.1%. YOLOv5 exhibited excellence with an AP of 96.8%, recall of 66.1%, and IoU of 74.0%. Conversely, SSD displayed a comparatively lower AP of 59.4%, recall of 38.6%, and IoU of 49.2%. Performance metrics such as AP, recall, and IoU showed variations for Faster RCNN and CenterNet.
@@
Figure 14: Inference examples of YOLOv8, DETR, SSD, and CenterNet on unseen images from the Google Earth HRPlanesv2 dataset and Airbus GDIT. Green boxes showcase the accurate detections, ‘FP’ stands for "false positive", ‘ND’ for "no detection" and ‘IE’ for "inaccurate estimation".
Figure 15: Inference examples of YOLOv5, RTMDet, RetinaNet and Faster-RCNN on another set of unseen images from Google Earth HRPlanesv2 dataset and Airbus GDIT. Green boxes showcase the accurate detections, ‘FP’ stands for "false positive", ‘ND’ for "no detection".
@@
Table 5 summarizes the object detection performance metrics of various models on remote sensing images across different subsets—train, test, validation, and all the datasets. Furthermore, as expected, the YOLO models stand out as the top real-time prediction models, with YOLOv10 being the best with 54.3 frames per second (FPS) due to its reduced number of trainable parameters. The overall performances showed that YOLOv5 emerged as the top-performing algorithm across all subsets in terms of accuracy, with the highest evaluation metrics AP and IoU. Notably, YOLOv5 demonstrated commendable recall rates and IoU scores, positioning it as the leading algorithm for aircraft detection in this study. Additionally, YOLOv5 also yielded a real-time performance of 37.4 FPS, usually acceptable in industrial environments. Conversely, SSD consistently exhibited comparatively lower performance metrics, indicating challenges in accurately detecting aircraft instances across subsets. While other algorithms, including YOLOv8, Faster RCNN, and CenterNet, displayed varying degrees of success, YOLOv5 consistently outperformed them in terms of AP, recall, and IoU.
Figure 16: Estimated evaluation metrics when inferencing the 8 models (initially trained on HRPlanesV2) on all images from unseen subsets ‘Train’ (a), ‘Test’ (b) and ‘Validation’ (c) from the GDIT aircraft dataset.
Table 5: Trained models for aircraft detection and classification performance on remote sensing images. Tr. subset refers to all images from ‘train’, Te. subset refers to all images from ‘test’, Val. subset refers to all images from ‘validation’, and All dataset refers to all images in the dataset.
Architecture Real-Time Perf. Metric Tr. Subset Te. Subset Val. Subset All Dataset
YOLOv8 43.4 FPS AP 0.919 0.903 0.900 0.907
@@ 6 Conclusion
This study presents a comprehensive evaluation of aircraft detection algorithms in satellite imagery, namely YOLO (v5 and v8), Faster RCNN, CenterNet, RetinaNet, RTMDet and DETR, specifically focusing on the HRPlanesV2 dataset and extending the assessment to subsets of the GDIT dataset: train, test, and validation. The used training setup involves training the aforementioned object detection algorithms on the dataset for 500 epochs on three NVIDIA RTX A6000 GPUs, with each GPU having 48 GB of memory. The results of the evaluation demonstrate the adaptability and robustness of the trained object detection algorithms. Among these algorithms, YOLOv5 emerges as the standout performer, achieving the highest mean average precision (mAP) of 0.99, highlighting its precision and robustness. YOLOv8 and v10 closely follow, further emphasizing the effectiveness of the YOLO architecture in aerial object detection. On the other side, SSD displayed the lowest performances in both the training and evaluation. Furthermore, the evaluation extends to the GDIT dataset, providing a more comprehensive assessment by deploying the trained network on other scenarios, including different satellite imagery sources. By employing evaluation metrics such as average precision, recall, and intersection over union, YOLOv5 still consistently outperforms the other algorithms, demonstrating superior performance across all subsets. This solidifies YOLOv5’s position as the top-performing algorithm for aircraft detection, characterized by obtained AP, recall, and IoU scores.
The topic of this research offers an extensive overview and comparative study, providing an in-depth analysis of the performance, accuracy, and computational demands of object detection methods. The insights gained from this study significantly enhance the decision-making process for selecting the most effective aircraft localization techniques in satellite imagery, supported by detailed training and validation processes using the HRPlanesv2 dataset and additional validation with the GDIT dataset.